# Improving Robustness Estimates in Natural Language Explainable AI though Synonymity Weighted Similarity Measures

### v1

@@ Abstract
Explainable AI (XAI) has seen a surge in recent interest with the proliferation of powerful but intractable black-box models. Moreover, XAI has come under fire for techniques that may not offer reliable explanations. As many of the methods in XAI are themselves models, adversarial examples have been prominent in the literature surrounding the effectiveness of XAI, with the objective of these examples being to alter the explanation while maintaining the output of the original model. For explanations in natural language, it is natural to use measures found in the domain of information retrieval for use with ranked lists to guide the adversarial XAI process. We show that the standard implementation of these measures are poorly suited for the comparison of explanations in adversarial XAI and amend them by using information that is discarded, the synonymity of perturbed words. This synonymity weighting produces more accurate estimates of the actual weakness of XAI methods to adversarial examples.
@@ 1 Introduction
Intractably complex models have proliferated in recent years as a result of their efficacy over their simpler relatives. And while efficacy is a necessary component of a good model, it is difficult to trust a black-box and so the question of “How does this model actually work?” follows. The lack of answers to this question has slowed the adoption of powerful models in fields where flawed models can have severe consequences, such as medicine Markus et al. (2021); Tjoa and Guan (2020).
@@
The discipline of Explainable AI (XAI) attempts to answer our question by generating an explanation of how the model generates its output. However, the XAI process that generates the explanation is itself a model, and while by design the XAI model can be inherently understood, we return to the fundamental question of any model “How effective is it?” Unfortunately, this question is nearly as difficult as our original query. As the models we attempt to explain are themselves incapable of being (comprehensively) understood, we have no ground truth against which to compare the output of the XAI method. However, we can judge the quality of an XAI method using criteria that avoid an understanding of the original intractable model and instead focus on the consistency of the output of the XAI method when subject to perturbations that retain the meaning of the original input. The property is called stability, or robustness, and is a necessary attribute required to trust an explanation. In general, stability is the intuitive property where insignificant changes to the input lead to only small changes to the output. An XAI process that lacks stability produces substantially different explanations under similar inputs and so is of limited, if any, use.
@@
To assess stability, it is necessary to compare the output of the explainability method, and in doing so, some measure is required to provide the comparisons. These similarity measures are the engine that drives the adversarial search process and so suboptimal measures will result in incorrect conclusions of an XAI methods stability. The similarity measures here are subject to two primary flaws. The first is the sensitivity, where two explanations are fundamentally similar, but minor differences are drastically magnified, resulting in effective false positive attack success. The second is the opposite, indifference or coarseness where important differences between the explanations are not captured. Although the measures and metrics in XAI have been directly discussed in previous research Hoffman et al. (2019), the focus was primarily on evaluating the quality of the explanations rather than the quality of the comparison between the explanations. In particular, the work Burger et al. (2023) established desiderata in text-based adversarial XAI using common desirable attributes in information retrieval Kumar and Vassilvitskii (2010), and used these to motivate an algorithm that mitigates the previously mentioned flaws. However, no prior work on text-based XAI has built purpose-driven measures and instead relied on standard formulations.
@@
Table 1 provides an example of how standard measures tend to be overly sensitive despite the explanations being intuitively very similar. We see large increases in similarity under a measure that uses synonymity weighting on perturbed features, which follows our intuition, as the words fulfill the same structure and possess much of the same meaning. Our focus on altering the similarity measures also has the advantage of reduced complexity in that the search process does not need to be taken into account in the attack algorithm. Since existing measures tend towards the sensitive side, careful choice in how which words are selected for perturbation is required. Without this limitation, we can now freely replace the search process, which also gives us the ability to use many adversarial natural language processes directly which otherwise would require modifications such as those in Burger et al. (2023).
Our contributions are: (1) The extension of some common measures in information retrieval to use synonymity weighting, focused on providing a more appropriate choice for adversarial methods in XAI. We leverage prior works in element-weighted similarity to allow a more accurate comparison between explanations which provides for a superior understanding of an XAI method’s robustness. (2) A comprehensive comparison against adversarial examples previously determined to be successful at some threshold of similarity under the standard measures. We show that prior conclusions about XAI instability drawn using certain existing measures may be inaccurate, especially for methods with naive search procedures.
@@ 2 Background and Related Work
Prior work on XAI stability has focused on evaluating models using tabular or image data in various interpretation methods, which often use small perturbations of the input data to generate appreciably different explanations Alvarez-Melis and Jaakkola (2018); Ghorbani et al. (2019); Alvarez-Melis and Jaakkola (2018), or generate explanations consisting of arbitrary features Slack et al. (2020). As our goal is to leverage the idea of synonymity inherent to features of an XAI explanation, we concentrate exclusively on XAI methods that operate on text-based input. Previous work directly involves adversarial perturbations for text-based XAI using a variety of search processes and similarity measures / distance metrics Sinha et al. (2021); Ivankay et al. (2022); Burger et al. (2023). To simplify further discussion, we refer to any applicable similarity measure or distance metric as simply a similarity measure.
Popular post hoc XAI methods like Lime Ribeiro et al. (2016) provide a ranked list of features ordered by their importance with respect to the desired model output, and comparisons between ranked lists is a standard task in the field of information retrieval. In particular, previous work has extended commonly used measures for ranked lists to allow weighted similarity of elements Kumar and Vassilvitskii (2010); Sculley (2007). However, it is important to note that the use of these specific measures is not required to apply synonymity-weighted similarity.
@@ 2.1 Adversarial Attack Process
Let $\mathbf{M}$ be a model that can be explained using the XAI method $\mathbf{E}$. Let $\mathbf{I}$ be some input to $\mathbf{M}$ we desire the explanation to be centered upon, that is we desire to understand how $\mathbf{M}$ uses the components of $\mathbf{I}$ to produce some output. $\mathbf{M(I)}$ is fed into $\mathbf{E}$ to generate an initial explanation $\mathbf{A}$ used as the basis of comparison against future perturbed versions of $\mathbf{I}$, termed $\mathbf{B}$. The input $\mathbf{I}$ is then perturbed by replacing some subset, usually a single word, with an appropriate substitution. This substitution is a close synonym, as judged with respect to a given embedding space, and is subject to certain constraints such as the grammatical category or location in the unperturbed explanation $\mathbf{I}$. The goal here is to generate a replacement while maintaining the fundamental meaning and structure of $\mathbf{I}$ despite the constituent words differing. The perturbed explanation $\mathbf{B}$ is compared to $\mathbf{I}$ using some similarity measure, and the process repeats with $\mathbf{B}$ as the new document to be perturbed until some termination condition. We note that prior work in adversarial XAI has enforced a single word to single word replacement for the perturbations due to the extensive reduction in search complexity and thus computation time (which remains a significant bottleneck even with highly greedy search processes). This restriction to a single word replacement scheme imposes an implicit mapping which can then be used to implement synonymity weighted similarity.
@@ 3 Mappings Between Explanations
Creating this mapping is simple if $|\mathbf{A}|=|\mathbf{B}|$ and the perturbation process is restricted to single-word substitutions. Any element $a\in\mathbf{A}$ located at index $i$ in that is perturbed to some value $b$ must either be located at some index $j$ in $\mathbf{B}$ or no longer part of the surrogate model and so is missing from explanation $\mathbf{B}$. For the latter case, this generally means eliminated in importance. However, if the search process does not exclude highly ranked features, these can be selected (and often are without careful choice of search constraints), resulting in substantial differences under certain similarity measures despite little substantial difference in meaning. This is one of the major issues that we seek to address with synonymity weighting.
@@
Some measures can be simply extended to handle disjoint elements. For example, Kendall’s Tau (Rank Distance) which counts the number of dissonant pairs between two ranked lists, can be easily extended to the comparison of different sized lists simply by declaring whatever excess element(s) that exist in the larger list to be automatically dissonant. But even for measures whose structure discards element pairings (generally set intersection-based), we must establish a mapping to apply the synonymity weighting. We will assume that all elements contained within the original explanation $\mathbf{A}$ are mapped, either to an element in $\mathbf{B}$ or the null mapping which indicates that the measure specific penalty should be applied. Additionally, multi-word substitutions for the perturbation method increases the difficulty of maintaining appropriate syntactic consistency and have not seen much use in adversarial explanations in XAI. As such, we will assume that all perturbation methods replace at most one word per iteration.
@@ 4 Constructing Weighted Similarity
For the interval estimate, we define a function $Syn(a,b)\longrightarrow[0,1]$ where $a,b$ are features within explanations $\mathbf{A},\mathbf{B}$ respectively. The function $Syn(\cdot)$ returns a value proportional to the synonymity between the features $a$ and $b$. We constrain the definition of $Syn(a,b)$ minimally, with the only condition required being $Syn(x,x)=1$, where $1$ is the absolute maximum similarity possible. In particular, the interval itself is subject to alteration. We choose $[0,1]$ to provide a simple representation for the proportion of similarity between two words and so make the analysis easier to perform, but this is not required and other choices are possible. Of these alternative choices, the interval $[-1,1]$ may be the most intuitive choice by letting $-1$ indicate that $b$ is an antonym of $a$, and 00 being that $b$ is completely unrelated to $a$.
@@
For example, let $a=good$, $b=bad$, and $c=frog$. Now for the interval $[0,1]$, $Syn(a,b)=Syn(a,c)=0$ as neither word is synonymous with good. For the interval $[-1,1]$ $Syn(a,b)\approx-1$ and $Syn(a,c)=0$ as bad is an antonym of the adjective good. Since part of speech checking is a commonly imposed constraint in text-based adversarial XAI, we assume that there is no ambiguity between words that span multiple syntactic categories such as good (noun) and good (adjective) should the synonymity measure be capable of handling this distinction. In particular, measures that incorporate embedding vectors do not necessarily make a distinction between identical words with multiple meanings (the embeddings are not multisense). These embeddings also will often not produce values with such obvious delineation. Good and frog will likely possess some similarity $>0$ by construction of the embedding. Choosing an optimal embedding is likely task specific and outside the scope of this discussion.
@@
We can also represent synonymity with a method akin to that of a traditional thesaurus, that is, there is a fixed collection of synonyms for a given word with any other word by definition not being synonymous. For a given word $a$, define the set of valid synonyms for $a$ as $\mathbf{S}_{a}$. Then our similarity function $Syn(a,b)$ is the characteristic function $\chi({\mathbf{S}_{a}}(b))$ where if $b\in\mathbf{S}_{a}$ then $Syn(a,b)=1$, otherwise it is 00. This “thesaurus” method may prove appealing to those who are dubious of the quality of estimates produced by comparison between word embeddings. This strategy is particularly useful for adversarial XAI methods whose search process focuses on perturbing important features first, as the change in similarity between adversarial substitutions can be unreasonably high for methods when large weights are applied to the top subset of features.
@@
While both possibilities for synonymity weighting are reasonable, each comes with some amount of subjectivity. For the interval representation, the continuous-valued output retains some measure of comparison between synonyms, as not all synonyms may be truly identical substitutions. However, the question of how closely related are the words good and great is subject to personal interpretation. Of course, words will generally be represented in terms of some embedding, but different embeddings may have different representations of a word and subsequently different calculations for the resulting similarity. Defining a set of valid synonyms avoids the subjectivity associated with the numerical representation of the synonym, but the question now shifts as to why certain elements are contained within the set. The remainder of the paper uses the interval-based representation that is calculated using embedding vectors and restricted to $[0,1]$ exclusively as the given method used to construct and test synonymity weighting to simplify the exposition and experimental verification.
@@ 5.1 Similarity Measures
We extend Kendall’s Tau to use synonymity weighting by adjusting the value of a mapped dissonant pair $a,b$ at an equal location by multiplying by $1-Syn(a,b)$. For highly synonymous replacements ($Syn(a,b)\to 1$) this assigns a distance close to zero for the mapped pair, and for dissimilar words it approaches the default distance.
@@
(3) Spearman’s footrule (Eq. 6) is the sum of the difference between the location $i$ of each feature $a\in\mathbf{A}$ to its corresponding location $j$ in $\mathbf{B}$. Spearman’s footrule is effectively the $L_{1}$ distance between ranked lists. Like Kendall’s Tau, the footrule is by default not intended for disjoint lists but can be altered by applying a penalty $p$ for disjoint elements. Here we use the formulation with a location parameter from Fagin et al. (2002) which is designed for top k lists and choose a penalty value of $\frac{k}{2}$.
@@
We extend the footrule to use synonymity weighting by taking the summation of three mutually exclusive conditions that a pair of features within the explanations may have. For features unchanged in both explanations, we calculate the distance as normal. For features $a,b$ with a mapping $a\to b$ we calculate the minimum between the standard distance divided by the synonymity between features $Syn(a,b)$ and the maximum possible distance under the default footrule $|A|-1$.
@@
(4) Rank-biased Overlap (RBO) is a sum of successively larger intersections, each weighted by a term in a convergent series. This weighting scheme is controlled by a parameter $p\in(0,1)$ that can be adjusted to ascribe more or less weight to the top k features. In general, values further down the list are weighted as less significant, which is often the case in XAI, as only the top few features are of interest to many end users Verma et al. (2020). RBO Webber et al. (2010) is defined in Eq. 8 where $d$ is the current depth of the ranking, and $k$ is the maximum depth.

### v2

# Quantifying True Robustness: Synonymity-Weighted Similarity for Trustworthy XAI Evaluation

@@ Abstract
Adversarial attacks challenge the reliability of Explainable AI (XAI) by altering explanations while the model’s output remains unchanged. The success of these attacks on text-based XAI is often judged using standard information retrieval metrics. We argue these measures are poorly suited in the evaluation of trustworthiness, as they treat all word perturbations equally while ignoring synonymity, which can misrepresent an attack’s true impact. To address this, we apply synonymity weighting, a method that amends these measures by incorporating the semantic similarity of perturbed words. This produces more accurate vulnerability assessments and provides an important tool for assessing the robustness of AI systems. Our approach prevents the overestimation of attack success, leading to a more faithful understanding of an XAI system’s true resilience against adversarial manipulation.
@@ 1 Introduction
Intractably complex models have proliferated in recent years as a result of their efficacy over their simpler relatives. And while efficacy is a necessary component of a good model, it is difficult to trust a black-box and so the question of “How does this model actually work?” follows. The lack of answers to this question has slowed the adoption of powerful models in fields where flawed models can have severe consequences, such as medicine [markus2021role, tjoa2020survey].
@@
The discipline of Explainable AI (XAI) attempts to answer our question by generating an explanation of how the model generates its output. However, the XAI process that generates the explanation is itself a model, and while by design the XAI model can be inherently understood, we return to the fundamental question of any model “How effective is it?” Unfortunately, this question is nearly as difficult as our original query. As the models we attempt to explain are themselves incapable of being (comprehensively) understood, we have no ground truth against which to compare the output of the XAI method.
However, we can judge the quality of an XAI method using criteria that avoid an understanding of the original intractable model and instead focus on the consistency of the output of the XAI method when subject to perturbations that retain the meaning of the original input. The property is called stability, or robustness, and is a necessary attribute required to trust an explanation. In general, stability is the intuitive property where insignificant changes to the input lead to only small changes to the output. An XAI process that lacks stability produces substantially different explanations under similar inputs and so is of limited, if any, use.
@@
To assess stability, it is necessary to compare the output of the explainability method, and in doing so, some measure is required to provide the comparisons. These similarity measures are the engine that drives the adversarial search process and so suboptimal measures will result in incorrect conclusions of an XAI methods stability. The similarity measures here are subject to two primary flaws. The first is the sensitivity, where two explanations are fundamentally similar, but minor differences are drastically magnified, resulting in effective false positive attack success. The second is the opposite, indifference or coarseness where important differences between the explanations are not captured. Although the measures and metrics in XAI have been directly discussed in previous research [hoffman2019metrics], the focus was primarily on evaluating the quality of the explanations rather than the quality of the comparison between the explanations. In particular, the work [burger2023explanations] established desiderata in text-based adversarial XAI using common desirable attributes in information retrieval [GenDistKumar], and used these to motivate an algorithm that mitigates the previously mentioned flaws. However, no prior work on text-based XAI has built purpose-driven measures and instead relied on standard formulations.
@@
Table 1 provides an example of how standard measures tend to be overly sensitive despite the explanations being intuitively very similar. We see large increases in similarity under a measure that uses synonymity weighting on perturbed features, which follows our intuition, as the words fulfill the same structure and possess much of the same meaning. Our focus on altering the similarity measures also has the advantage of reduced complexity in that the search process does not need to be taken into account in the attack algorithm. Since existing measures tend towards the sensitive side, careful choice in which words are selected for perturbation is required. Without this limitation, we can now freely replace the search process, which also gives us the ability to use many adversarial natural language processes directly which otherwise would require modifications such as those in [burger2023explanations].
Our contributions are: (1) The extension of some common measures in information retrieval to use synonymity weighting, focused on providing a more appropriate choice for adversarial methods in XAI. We leverage prior works in element-weighted similarity to allow a more accurate comparison between explanations which provides for a superior understanding of an XAI method’s robustness. (2) A comprehensive comparison against adversarial examples previously determined to be successful at some threshold of similarity under the standard measures. We show that prior conclusions about XAI instability drawn using certain existing measures may be inaccurate, especially for methods with naive search procedures. Ultimately, without a reliable way to measure the impact of adversarial perturbations, we cannot establish the trustworthiness of an explanation, which is a critical prerequisite for AI Safety in high-stakes domains.
@@ 2 Background and Related Work
Prior work on XAI stability has focused on evaluating models using tabular or image data in various interpretation methods, which often use small perturbations of the input data to generate appreciably different explanations [alvarezmelis2018robustness, InterpretationNN, alvarezmelis2018robustness], or generate explanations consisting of arbitrary features [slack2020fooling]. As our goal is to leverage the idea of synonymity inherent to features of an XAI explanation, we concentrate exclusively on XAI methods that operate on text-based input. Previous work directly involves adversarial perturbations for text-based XAI using a variety of search processes and similarity measures / distance metrics [sinha2021perturbing, ivankay2022fooling, burger2023explanations, burger2025]. To simplify further discussion, we refer to any applicable similarity measure or distance metric as simply a similarity measure.
Popular post hoc XAI methods like Lime [LIME_Ribeiro] provide a ranked list of features ordered by their importance with respect to the desired model output, and comparisons between ranked lists is a standard task in the field of information retrieval. In particular, previous work has extended commonly used measures for ranked lists to allow weighted similarity of elements [GenDistKumar, Sculley2007RankAF]. However, it is important to note that the use of these specific measures is not required to apply synonymity-weighted similarity.
@@ 2.1 Adversarial Attack Process
Let $\mathbf{M}$ be a model that can be explained using the XAI method $\mathbf{E}$. Let $\mathbf{I}$ be some input to $\mathbf{M}$ we desire the explanation to be centered upon, that is we desire to understand how $\mathbf{M}$ uses the components of $\mathbf{I}$ to produce some output. $\mathbf{M(I)}$ is fed into $\mathbf{E}$ to generate an initial explanation $\mathbf{A}$ used as the basis of comparison against future perturbed versions of $\mathbf{I}$, termed $\mathbf{B}$.
The input $\mathbf{I}$ is then perturbed by replacing some subset, usually a single word, with an appropriate substitution. This substitution is a close synonym, as judged with respect to some measure, and is subject to certain constraints such as the grammatical category or location in the unperturbed explanation $\mathbf{I}$. The goal here is to generate a replacement while maintaining the fundamental meaning and structure of $\mathbf{I}$. The perturbed explanation $\mathbf{B}$ is compared to $\mathbf{I}$ using some similarity measure, and the process repeats with $\mathbf{B}$ as the new document to be perturbed until some termination condition.
We note that prior work in adversarial XAI has enforced a single word to single word replacement for the perturbations due to the extensive reduction in search complexity and thus computation time (which remains a significant bottleneck) This restriction to a single word replacement scheme imposes an implicit mapping which can then be used to implement synonymity weighted similarity.
@@ 3 Mappings Between Explanations
Creating this mapping is simple if $|\mathbf{A}|=|\mathbf{B}|$ and the perturbation process is restricted to single-word substitutions. Any element $a\in\mathbf{A}$ located at index $i$ in $\mathbf{A}$ that is perturbed to some value $b$ must either be located at some index $j$ in $\mathbf{B}$ or no longer part of the surrogate model and so is missing from explanation $\mathbf{B}$. For the latter case, this generally means eliminated in importance. However, if the search process does not exclude highly ranked features, these can be selected (and often are without careful choice of search constraints), resulting in substantial differences under certain similarity measures despite little substantial difference in meaning. This is one of the major issues that we seek to address with synonymity weighting.
@@
Some measures can be simply extended to handle disjoint elements. For example, Kendall’s Tau (Rank Distance) which counts the number of dissonant pairs between two ranked lists, can be easily extended to the comparison of different sized lists simply by declaring whatever excess element(s) that exist in the larger list to be automatically dissonant. But even for measures whose structure discards element pairings (generally set intersection-based), we must establish a mapping to apply the synonymity weighting.
We will assume that all elements contained within the original explanation $\mathbf{A}$ are mapped, either to an element in $\mathbf{B}$ or the null mapping which indicates that the measure specific penalty should be applied. Additionally, multi-word substitutions for the perturbation method increases the difficulty of maintaining appropriate syntactic consistency and have not seen much use in adversarial explanations in XAI. As such, we will assume that all perturbation methods replace at most one word per iteration.
@@ 4 Constructing Weighted Similarity
For the interval estimate, we define a function $Syn(a,b)\longrightarrow[0,1]$ where $a,b$ are features within explanations $\mathbf{A},\mathbf{B}$ respectively. The function $Syn(\cdot)$ returns a value proportional to the synonymity between the features $a$ and $b$. We constrain the definition of $Syn(a,b)$ minimally, with the only condition required being $Syn(x,x)=1$, where $1$ is the absolute maximum similarity possible. In particular, the interval itself is subject to alteration. We choose $[0,1]$ to provide a simple representation for the proportion of similarity between two words and so make the analysis easier to perform, but this is not required and other choices are possible. Of these alternative choices, the interval [1,1][-1,1] may be the most intuitive choice by letting 1-1 indicate that $b$ is an antonym of $a$, and 0 being that $b$ is completely unrelated to $a$.
@@
For example, let $a=good$, $b=bad$, and $c=frog$. Now for the interval $[0,1]$, $Syn(a,b)=Syn(a,c)=0$ as neither word is synonymous with good. For the interval [1,1][-1,1] $Syn(a,b)\approx-1$ and $Syn(a,c)=0$ as bad is an antonym of the adjective good. Since part of speech checking is a commonly imposed constraint in text-based adversarial XAI, we assume that there is no ambiguity between words that span multiple syntactic categories such as good (noun) and good (adjective) should the synonymity measure be capable of handling this distinction. In particular, measures that incorporate embedding vectors do not necessarily make a distinction between identical words with multiple meanings (the embeddings are not multisense). These embeddings also will often not produce values with such obvious delineation. Good and frog will likely possess some similarity >0>0 by construction of the embedding. Choosing an optimal embedding is likely task specific and outside the scope of this discussion.
@@
We can also represent synonymity with a method akin to that of a traditional thesaurus, that is, there is a fixed collection of synonyms for a given word with any other word by definition not being synonymous. For a given word $a$, define the set of valid synonyms for $a$ as $\mathbf{S}_{a}$. Then our similarity function $Syn(a,b)$ is the characteristic function $\chi({\mathbf{S}_{a}}(b))$ where if $b\in\mathbf{S}_{a}$ then $Syn(a,b)=1$, otherwise it is 0. This “thesaurus” method may prove appealing to those who are dubious of the quality of estimates produced by comparison between word embeddings. This strategy is particularly useful for adversarial XAI methods whose search process focuses on perturbing important features first, as the change in similarity between adversarial substitutions can be unreasonably high for methods when large weights are applied to the top subset of features.
@@
While both possibilities for synonymity weighting are reasonable, each comes with some amount of subjectivity. For the interval representation, the continuous-valued output retains some measure of comparison between synonyms, as not all synonyms may be truly identical substitutions. However, the question of how closely related are the words good and great is subject to personal interpretation. Of course, words will generally be represented in terms of some embedding, but different embeddings may have different representations of a word and subsequently different calculations for the resulting similarity. Defining a set of valid synonyms avoids the subjectivity associated with the numerical representation of the synonym, but the question now shifts as to why certain elements are contained within the set.
For our main experiments, we selected the widely-used GloVe-Twitter-25 model to generate similarity scores. The results of this primary analysis are detailed in the subsequent sections (and seen Tables 3-4, Figures 1-2). Then, to test the robustness of our conclusions to choice of embedding we conduct a sensitivity analysis, replicating our experiments with two methodologically different options: fastText, which leverages subword information, and a thesaurus-based method derived from WordNet (as described above). As our results will show, this analysis demonstrates that the core findings of our work are consistent across all three measures, reinforcing the general applicability of our proposed weighting method.
@@ 5.1 Similarity Measures
We extend Kendall’s Tau to use synonymity weighting by adjusting the value of a mapped dissonant pair $a,b$ at an equal location by multiplying by 1Syn(a,b)1-Syn(a,b). For highly synonymous replacements ($Syn(a,b)\to 1$) this assigns a distance close to zero for the mapped pair, and for dissimilar words it approaches the default distance.
@@
(3) Spearman’s footrule (Eq. 6) is the sum of the difference between the location $i$ of each feature $a\in\mathbf{A}$ to its corresponding location $j$ in $\mathbf{B}$. Spearman’s footrule is effectively the $L_{1}$ distance between ranked lists. Like Kendall’s Tau, the footrule is by default not intended for disjoint lists but can be altered by applying a penalty $p$ for disjoint elements. Here we use the formulation with a location parameter from [topklists] which is designed for top k lists and choose a penalty value of $\frac{k}{2}$.
@@
We extend the footrule to use synonymity weighting by taking the summation of three mutually exclusive conditions that a pair of features within the explanations may have. For features unchanged in both explanations, we calculate the distance as normal. For features $a,b$ with a mapping $a\to b$ we calculate the minimum between the standard distance divided by the synonymity between features $Syn(a,b)$ and the maximum possible distance under the default footrule |A|1|A|-1.
@@
(4) Rank-biased Overlap (RBO) is a sum of successively larger intersections, each weighted by a term in a convergent series. This weighting scheme is controlled by a parameter $p\in(0,1)$ that can be adjusted to ascribe more or less weight to the top k features. In general, values further down the list are weighted as less significant, which is often the case in XAI, as only the top few features are of interest to many end users [verma2020counterfactual]. RBO [10.1145/1852102.1852106] is defined in Eq. 8 where $d$ is the current depth of the ranking, and $k$ is the maximum depth.